# EmoSphere++: Emotion-Controllable Zero-Shot Text-to-Speech via Emotion-Adaptive Spherical Vector

### v1

@@ Abstract
Emotional text-to-speech (TTS) technology has achieved significant progress in recent years; however, challenges remain owing to the inherent complexity of emotions and limitations of the available emotional speech datasets and models. Previous studies typically relied on limited emotional speech datasets or required extensive manual annotations, restricting their ability to generalize across different speakers and emotional styles. In this paper, we present EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. We introduce a novel emotion-adaptive spherical vector that models emotional style and intensity without human annotation. Moreover, we propose a multi-level style encoder that can ensure effective generalization for both seen and unseen speakers. We also introduce additional loss functions to enhance the emotion transfer performance for zero-shot scenarios. We employ a conditional flow matching-based decoder to achieve high-quality and expressive emotional TTS in a few sampling steps. Experimental results demonstrate the effectiveness of the proposed framework.
@@ I Introduction
Emotions are interrelated in a highly systematic fashion [1]. For example, the emotion of sadness can be expressed with derivative states of primary emotions such as feeling hurt or lonely, depending on the style and intensity. Recently, emotional text-to-speech (TTS) technology has experienced rapid developments, increasing the interest in high-level interpretable emotion control [2, 3, 4, 5]. Controllable emotional TTS represents a breakthrough in reproducing human-like emotions in speech synthesis, thus enabling more emotionally intelligent interactions between humans and computers. Although researchers have made significant progress in controlling emotional intensity, the ability to precisely control emotional style must be further improved.
The modeling of diverse emotional styles and intensities represents a significant challenge in controllable emotion TTS. A reason for this challenge is that emotion style and intensity are even more subjective and complex than discrete emotion categories, making them difficult to model. Two general approaches for achieving controllable emotional TTS involve controlling conditioning features or manipulating internal emotion representations. That is, one approach uses conditioning features of emotion intensity, such as relative ranking matrices [6, 7, 8, 9, 10, 11], distance-based quantization [12], or voiced, unvoiced, and silence (VUS) states [13]. The alternative approach involves the manipulation of internal emotion representations through the application of scaling factors [14, 15] or interpolation of the embedding space [16]. However, despite these methods, the explicit control of emotion style and intensity remains a largely unexplored topic in emotional speech synthesis.
@@
Another approach to controlling emotional expression involves utilizing emotional dimensions. Compared to the discrete emotion approach, the dimensional approach, such as Russell’s circumplex model, provides a more precise method for capturing the nuances between different emotional states [1]. Recently, studies on TTS systems have attempted to control emotional attributes through the emotion dimension [17, 18]. In one of these studies, a prosody control block is extended by incorporating the continuous space of arousal and valence to allow interpretable emotional prosody control [18]. Another study proposes an expressive TTS model with a semi-supervised latent variable to control emotions in six discrete emotional states of arousal-valence combinations [17]. However, this setup requires labor-intensive annotations, which are more expensive to obtain than categorical labels and more susceptible to annotator bias. The emotional dimension model also exhibits limitations when explicitly controlling emotion style and intensity.
@@
Most emotional TTS systems utilize Sequence-to-sequence (Seq2Seq) models, which not only predict the duration of speech automatically but also learn feature mapping and alignment simultaneously [19, 20]. The attention mechanism in these models allows them to focus on the emotionally emphasized parts of an utterance [21]. However, Seq2Seq models face the typical challenges of auto-regressive models, such as long-term dependence and repetition problems. Furthermore, most emotional speech syntheses adopt fine-tuning to control emotion intensity on a single speaker dataset; however, some of these methods exhibit noticeably degraded speech quality [7, 10]. Researchers have explored acoustic models for emotion transfer to ensure the capture of expressiveness when synthesizing acoustic features. Existing methods focus on emotion transfer using discrete emotion labels, which overlook the complexity of emotions conveyed in human speech [15, 22]. To address this issue, a TTS system with high generalization capability in zero-shot emotion transfer scenarios must be implemented [23, 24, 25, 26, 2].
Figure 2: Illustration of coordinate transformations: (a) Cartesian-to-spherical coordinate transformation and (b) Emotion-adaptive coordinate transformation.
In psychology, several studies [27, 28] have been conducted to explore the frameworks and methods for measuring complex emotions that arise from more primary emotional states. Moreover, TTS models are being challenged by the increasing demand for personalized speech generation, focusing on achieving high generalization capability and producing high-quality speech. However, the following primary research problems exist: (1) how can emotional style and intensity be defined and modeled as derivatives of primary emotions, (2) how can high-quality emotional style be controlled and transferred in zero-shot scenarios, and (3) how can the synthesized emotional speech be evaluated. In this article, we address these challenges, with our primary contributions listed as follows:
@@
  • We introduce an emotion-adaptive coordinate transformation that models an emotion-adaptive spherical vector (EASV), allowing for subtle control over emotion style and intensity.
  • We introduce a multi-level style encoder for high- and low-level style adaptations of custom utterances. Particularly, we introduce additional loss functions to enhance the emotion transfer performance for zero-shot scenarios.
  • We propose a novel objective evaluation of spherical vector angle similarity (SVAS) to assess the emotional expressiveness of the speech.
@@ II-A Characterization of Emotions
The processes of defining and expressing emotions have garnered significant interest in psychology [29, 30, 1, 31]. Emotion theorists divide emotion theory into discrete [32, 27] and dimensional models [1, 33]. Notably, discrete models only represent the most straightforward way, whereas dimensional models aim to offer a continuous and fine-grained description.
Emotion labels correspond very closely to the categories that we use in our daily lives. Paul Ekman [32] derived six primary emotions: happiness, anger, disgust, sadness, anxiety, and surprise based on universally recognized facial expressions. However, this approach overlooks the nuanced variations of emotions. Plutchik’s emotion wheel [27] proposes eight primary emotions and suggests that all other emotions arise as derivative states of the primary emotions. Furthermore, varying the intensity of primary emotions on the emotion wheel can produce the wide range of emotions that humans experience. Although people can explicitly express emotions, modeling the relationships between discrete emotional states presents a challenge.
Researchers have introduced dimensional models to computationally interpret the relationships between emotional states. Russell’s circumplex model [1] suggests a two-dimensional circular space that spans the independent and bipolar dimensions of arousal and valence. Building on this, researchers have attempted to extend the model to the third dimension of dominance to denote the location of emotion within this space [33]. In the valence-arousal space, intensity is often equated with arousal; moreover, Reisenzein demonstrated that using the angle and length of the vector in polar coordinates is the only possible option for interpreting the relationships between emotions [28, 34]. Recently, the speech processing domain has seen various studies on emotion recognition that leverage the emotional dimension [35, 36, 37, 38].
Despite these efforts in psychology, the current literature on speech synthesis is still insufficient in effectively modeling and controlling the subtle variations of emotions. Inspired by several psychological theories [27, 28], we believe that the dimensional model allows speech synthesis models to express derived emotions of primary emotions, as shown in Fig. 1 (a). This approach enables researchers to generate, control, and manipulate a wide range of emotions more easily in real-life applications.
@@ II-B Controllable Emotional Speech Synthesis
Recently, speech synthesis models have exhibited significant developments [39, 40, 41]; therefore, controllable emotional speech synthesis research is being aggressively pursued [4, 3, 5]. Recording and annotating emotional datasets covering multiple emotional styles is incredibly challenging. Researchers typically use the following controllable emotional speech synthesis methods to utilize general emotional datasets: 1) emotion label- and 2) reference-based approaches.
The emotion label-based approach aims to properly model conditioning input to reflect the complex nature of emotions. Researchers typically model emotion intensity using a learned ranking function [42], as employed in [6, 7, 10, 8, 9, 11]. The ranking function [42] seeks a ranking matrix based on the relationships between the low-level and different high-level emotional expressions using support vector machines. The model receives the emotional intensity of emotional samples as a conditioning input for training. However, this method tends to rely on emotion labels and introduces bias into training through separate stages. Most models that use ranking functions adopt fine-tuning to control emotion intensity on a single-speaker dataset; however, some of these methods have noticeably degraded speech quality [7, 10]. Moreover, certain research studies have utilized conditioning input such as distance-based quantization [12] and VUS states [13] to model emotional intensity. However, these methods are still limited to several predefined emotion labels and lack differentiation among samples within the same emotion label.
As emotional speech synthesis often lacks multiple emotional style labels, reference-based approaches are famous for using reference audio to transfer emotional styles. Several studies have controlled emotion intensity through operations on representative emotion embedding. The scaling factors approach [14, 15] reflects fine-grained emotion representation through multiplication. In addition to the scaling approach, the interpolation approach proposed by [16] controls emotion intensity through an inter-to-intra emotional distance ratio algorithm. However, the structure of the embedding space influences model performance and complicates the process of finding proper parameters for scaling or interpolation.
However, these frameworks can neither be explicitly tuned such as label-based approaches nor capture fine-grained emotion representations such as reference-based approaches. This study aims to fill this research gap by explicitly modeling EASV of emotion style and intensity variations.
@@ II-C Style and Emotion Transfer in Text-to-Speech
Researchers in the TTS community have studied style and emotion transfer for decades. Some studies introduced the concept of global style tokens (GSTs) [43]. One such derivative is Mellotron [44], an autoregressive multi-speaker TTS model based on Tacotron that utilizes GSTs. Li et al. [15] proposed a module for disentangling and transferring emotions across different speakers to achieve a desired emotional tone while maintaining the identity of the speaker. However, these methods exhibit limitations in capturing different styles and zero-shot scenarios owing to the restricted speaker lookup table and the insufficient performance of the disentangling modules.
Conversely, iEmoTTS [45] and YourTTS [46] used pre-trained speaker embedding for robust zero-shot performance. Moreover, GenerSpeech [47] proposed a multi-level style adapter to obtain different styles, including a global latent representation with speaker and emotion features. However, previous research has lacked methods to effectively process style and disentangle speech factors. Our work handles well-formed style processing through multi-level style encoders and incorporation of additional loss to exhibit strong generalization performance for zero-shot scenarios.
Figure 3: Training diagram of the EmoSphere++ framework. The framework consists of three main modules: the text encoder, the multi-level style encoder, and the CFM decoder. The right section illustrates the detailed structure of the multi-level style encoder, which extracts high-level speaker, high-level emotion, and low-level emotion to form a multi-level style embedding for emotional speech synthesis.
@@ III Proposed Method
This paper introduces EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. We considered the emotion wheel theory [27], which suggests that all other emotions are derived from the primary emotions, and the circumplex model [1, 34], which represents emotions through coordinate transformations based on intensity. Inspired by these psychological theories, we believe that derivative emotions can be characterized by the coordinate transformation in the valence-arousal-dominance (VAD) dimension, reflecting the style and intensity of the primary emotions, as shown in Fig. 1 (c). In a previous work, a spherical emotion vector was introduced to model the style and intensity of emotions [3]. Building on this, we propose a novel EASV, enabling more precise manual manipulation of the synthesized emotional style and intensity. This framework facilitates the training of each high-level style pre-trained encoder with disentangling modules, allowing us to effectively generalize across zero-shot scenarios. The details are described in the following subsections.

### v2

@@ Abstract
Emotional text-to-speech (TTS) has advanced significantly, but challenges persist due to the complexity of emotions and limitations in emotional speech datasets and models. A key issue with previous studies is the reliance on limited emotional speech datasets or extensive manual annotations, which restrict generalization across different speakers and emotional styles. To address this, we propose EmoSphere++, an emotion-controllable zero-shot TTS model capable of generating expressive speech with fine-grained control over emotional style and intensity—without requiring manual annotations. We introduce a novel emotion-adaptive spherical vector that effectively captures emotional style and intensity, along with a joint attribute style encoder that enhances generalization to both seen and unseen speakers. To further improve emotion transfer in zero-shot scenarios, we introduce an additional disentanglement method to enhance the style transfer performance for zero-shot scenarios. Through both objective and subjective evaluations, we demonstrate the benefits of the proposed model in emotion style and intensity modeling, as well as its effectiveness in enhancing emotional expressiveness across both seen and unseen speakers.
@@ I Introduction
Emotions are interrelated in a highly systematic fashion [1]. For example, the emotion of sadness can be expressed with derivative states of primary emotions such as feeling hurt or lonely, depending on the style and intensity. In speech synthesis, the ability to generate expressive and controllable emotional speech is essential for creating natural and effective human-computer interactions, as emotions are nuanced and can manifest in varying styles and intensities. Recently, emotional text-to-speech (TTS) technology has experienced rapid developments, increasing the interest in global interpretable emotion control [2, 3, 4, 5]. Controllable emotional TTS represents a breakthrough in reproducing human-like emotions in speech synthesis, thus enabling more emotionally intelligent interactions between humans and computers. Although researchers have made significant progress in controlling emotional intensity, the ability to precisely control emotional style remains a challenge.
Modeling diverse emotional styles and intensities is a major challenge in controllable emotional TTS. Unlike discrete emotion categories, emotional style and intensity are highly subjective and complex, making them difficult to accurately represent. Two general approaches for achieving controllable emotional TTS involve controlling conditioning features or manipulating internal emotion representations. That is, one approach uses conditioning features of emotion intensity, such as relative ranking matrices [6, 7, 8, 9, 10, 11], distance-based quantization [12], or voiced, unvoiced, and silence (VUS) states [13]. The alternative approach involves the manipulation of internal emotion representations through the application of scaling factors [14, 15] or interpolation of the embedding space [16]. However, despite these methods, the explicit control of emotion style and intensity remains a largely unexplored topic in emotional speech synthesis.
@@
Another approach to controlling emotional expression involves utilizing emotional dimensions. Compared to the discrete emotion approach, the dimensional approach, such as Russell’s circumplex model, provides a more precise method for capturing the nuances between different emotional states [1]. Recently, studies on TTS systems have attempted to control emotional attributes through the emotion dimension [17, 18]. In one of these studies, a prosody control block is extended by incorporating the continuous space of arousal and valence to allow interpretable emotional prosody control [18]. Another study proposes an expressive TTS model with a semi-supervised latent variable to control emotions in six discrete emotional states of arousal-valence combinations [17]. However, this setup requires labor-intensive annotations, which are more expensive to obtain than categorical labels and more susceptible to annotator bias. The emotional dimension model also exhibits limitations when explicitly controlling emotion style and intensity. To address these challenges, EmoSphere-TTS [3] models derivative emotions through emotional attribute prediction and discrete emotion labels, enabling explicit control over emotion style and intensity. However, limitations persist due to reliance on predefined emotion and speaker labels.
@@
Most emotional TTS systems utilize Sequence-to-sequence (Seq2Seq) models, which not only predict the duration of speech automatically but also learn feature mapping and alignment simultaneously [19, 20]. The attention mechanism in these models allows them to focus on the emotionally emphasized parts of an utterance [21]. However, Seq2Seq models face the typical challenges of auto-regressive models, such as long-term dependence and repetition problems. Furthermore, most emotional speech syntheses adopt fine-tuning to control emotion intensity on a single speaker dataset; however, some of these methods exhibit noticeably degraded speech quality [7, 10]. Researchers have explored acoustic models and additional discriminators for emotion transfer to enhance the capture of expressiveness when synthesizing acoustic features [22, 23, 3, 24]. However, existing methods primarily focus on emotion transfer using discrete emotion labels, which overlook the complexity of emotions conveyed in human speech [15, 25]. Furthermore, these approaches often rely on additional discriminators to enhance expressiveness, adding complexity to the model while struggling to fully capture emotional nuance. To address these challenges, a TTS system capable of generalizing across zero-shot emotion transfer scenarios is needed, ensuring more accurate emotion synthesis without relying on predefined labels [26, 27, 28, 29, 2].
As previously discussed, existing study and prior work [3] face the following challenges: (1) Defining and modeling emotional style and intensity as derivatives of primary emotions, while accounting for characteristics such as the distribution of emotion categories; (2) Effectively integrating global and fine-grained emotion representations to enhance emotional expressiveness, while ensuring robust generalization across unseen speakers and emotions; (3) Designing a TTS system capable of achieving high generalization and expressive capability in zero-shot style transfer scenarios, without relying on additional modules; and (4) Evaluating synthesized emotional speech beyond global emotion assessment, to include subjective measurement of detailed emotion styles. To build upon these discussions, we were inspired by psychological studies [30, 31] that have explored frameworks and methods for measuring complex emotions that arise from more primary emotional states. Additionally, with the increasing demand for personalized speech generation, we aimed to address the challenges in TTS models by focusing on achieving high generalization capability and producing high-quality speech. In this article, we present the following key contributions:
@@
  • We introduce an emotion-adaptive coordinate transformation that models the emotion-adaptive spherical vector (EASV), enabling more interpretable and controllable synthesis of emotion style and intensity.
  • We introduce a joint attribute style encoder along with an additional disentanglement module, enabling the model to perform emotion transfer even in zero-shot scenarios where reference speakers and emotions are not explicitly labeled.
  • We propose a novel objective evaluation method, spherical vector angle similarity (SVAS), to evaluate overall emotion accuracy while also capturing subtle variations in speech emotion styles with greater precision.
@@ II-A Characterization of Emotions
The processes of defining and expressing emotions have garnered significant interest in psychology [32, 33, 1, 34]. Emotion theorists divide emotion theory into discrete [35, 30] and dimensional models [1, 36]. Discrete models represent emotions as distinct, separate categories, while dimensional models provide a continuous and fine-grained description, capturing the complexity and variability of emotional experiences.
Emotion labels correspond very closely to the categories we use in our daily lives. Paul Ekman [35] derived six primary emotions: happiness, anger, disgust, sadness, anxiety, and surprise based on universally recognized facial expressions. However, this approach overlooks the nuanced variations of emotions. For instance, Plutchik’s emotion wheel [30] proposes eight primary emotions and suggests that all other emotions arise as derivative states of the primary emotions. By adjusting the intensity of primary emotions on the wheel, a broader spectrum of emotional experiences can be represented. Although people can explicitly express emotions, modeling the relationships between discrete emotional states presents a challenge.
Researchers have introduced dimensional models to computationally interpret the relationships between emotional states. Dimensional models represent emotions along three continuous dimensions [1, 36]: valence represents the positivity or negativity of an emotion, arousal indicates the intensity of the emotion provoked by a stimulus, and dominance denotes the level of control exerted by the stimulus. Russell’s circumplex model [1] suggests a two-dimensional circular space that spans the independent and bipolar dimensions of arousal and valence. Building on this, researchers have attempted to extend the model to the third dimension of dominance to denote the location of emotion within this space [36]. In the valence-arousal space, intensity is often equated with arousal; moreover, Reisenzein demonstrated that using the angle and length of the vector in polar coordinates is the only possible option for interpreting the relationships between emotions [31, 37]. Recently, the speech processing domain has seen various studies on emotion recognition that leverage the emotional dimension [38, 39, 40, 41].
Despite these efforts in psychology, the current literature on speech synthesis is still insufficient in effectively modeling and controlling the subtle variations of emotions. Inspired by several psychological theories [30, 31], we hypothesize that the dimensional model allows speech synthesis models to express derived emotions of primary emotions, as shown in Fig. 1 (a). This approach enables researchers to generate, control, and manipulate a wide range of emotions more easily in real-life applications.
Figure 2: Illustration of coordinate transformations: (a) Cartesian-to-spherical coordinate transformation [3] and (b) Emotion-adaptive coordinate transformation. In (a), the center $M$ represents the central coordinates of the neutral state, whereas (b) introduces $M_{k}$, which serves as the representative center that reflects the distributions of both neutral and target emotions.
@@ II-B Controllable Emotional Speech Synthesis
Recently, speech synthesis models have exhibited significant developments [23, 42, 43]; therefore, controllable emotional speech synthesis research is being aggressively pursued [4, 3, 5]. Researchers typically use the following controllable emotional speech synthesis methods to utilize general emotional datasets: 1) emotion label- and 2) reference-based approaches.
The emotion label-based approach aims to properly model conditioning input to reflect the complex nature of emotions. Researchers typically model emotion intensity using a learned ranking function [44], as employed in [6, 7, 10, 8, 9, 11]. The ranking function [44] seeks a ranking matrix based on the relationships between the dimensional-driven and different global emotional expressions using support vector machines. The model receives the emotional intensity of emotional samples as a conditioning input for training. However, this method tends to rely on emotion labels and introduces bias into training through separate stages. Most models that use ranking functions adopt fine-tuning to control emotion intensity on a single-speaker dataset; however, some of these methods have noticeably degraded speech quality [7, 10]. Moreover, certain research studies have utilized conditioning input such as distance-based quantization [12] and VUS states [13] to model emotional intensity. However, these methods are still limited to several predefined emotion labels and lack differentiation among samples within the same emotion label.
As emotional speech synthesis often lacks multiple emotional style labels, reference-based approaches are famous for using reference audio to transfer emotional styles. Several studies have controlled emotion intensity through operations on representative emotion embedding. The scaling factors approach [14, 15] reflects fine-grained emotion representation through multiplication. In addition to the scaling approach, the interpolation approach proposed by [16] controls emotion intensity through an inter-to-intra emotional distance ratio algorithm. Despite these techniques, the structure of the embedding space influences model performance and complicates the process of finding optimal parameters for scaling or interpolation.
However, these methods cannot be tuned explicitly like label-based approaches, nor do they capture the fine-grained emotion representations achievable by reference-based methods. EmoSphere-TTS [3] solves this by proposing a spherical emotion vector to control the emotional style and intensity of the synthetic speech. However, the lack of consideration for emotion category distribution in emotion style and intensity modeling can lead to unnatural variations in certain styles and intensities. This study addresses the lack of emotion category distribution-based modeling by explicitly modeling emotion style and intensity variations based on EmoSphere-TTS [3], thereby bridging this research gap.
@@ II-C Style and Emotion Transfer in Text-to-Speech
The TTS research community has long explored methods for style and emotion transfer. A significant advancement came with the introduction of global style tokens (GSTs) [45], which provided a framework for capturing and transferring speaking styles. One such derivative is Mellotron [46], an autoregressive multi-speaker TTS model based on Tacotron that utilizes GSTs. Li et al. [15] proposed a module for disentangling and transferring emotions across different speakers to achieve a desired emotional tone while maintaining the identity of the speaker. However, these methods exhibit limitations in capturing different styles and zero-shot scenarios owing to the restricted speaker lookup table and the insufficient performance of the disentangling modules.
In response, iEmoTTS [47] and YourTTS [48] used pre-trained speaker embedding for robust zero-shot performance. Additionally, GenerSpeech [49] proposed a multi-level style adapter to obtain different styles, including a global latent representation with speaker and emotion features. However, previous research has lacked methods to effectively process style and disentangle speech factors. Our work handles well-formed style processing through joint attribute style encoders and incorporation of additional loss to exhibit strong generalization performance for zero-shot scenarios.
II-D Speaker and Emotion Feature Disentanglement Methods
Some prosodic features are inherently associated with the speaker’s identity, making complete disentanglement challenging. Therefore, an effective disentangling method is critical to adequately separate speaker identity from emotion-related prosodic features, ensuring clear and accurate emotional transfer without compromising the target speaker’s timbre. Researchers typically implement the disentanglement module via explicit labels, such as the gradient reversal layer (GRL) [50], as employed in [47, 22]. However, using explicit labels for disentanglement introduces a trade-off between preserving emotional information and separating speaker identity, making hyperparameter optimization challenging and leading to suboptimal performance and synthesis quality [15]. Vector quantization (VQ) [51] offers an alternative approach to separating information without relying on explicit labels. Although VQ is effective in separating information without explicit labels [47, 49], it often results in unintended information loss and requires complex optimization to balance compression with reconstruction quality. To address these issues, [15] proposed an orthogonal loss to compensate for the embedding of emotion for the loss of emotional information caused by the disentanglement of speaker information [52]. Building on these advancements, we aim to minimize information loss during disentanglement using an orthogonal loss-based approach while enhancing the expressiveness of synthesized speech.
@@ III Proposed Method
This paper introduces EmoSphere++, an emotion-controllable zero-shot TTS model that can control emotional style and intensity to resemble natural human speech. Our work is grounded in the emotion wheel theory [30], which suggests that all other emotions are derived from the primary emotions, and the circumplex model [1, 37], representing emotions through coordinate transformations based on intensity. We propose that derivative emotions can be characterized by the coordinate transformation in the valence-arousal-dominance (VAD) dimension, reflecting the style and intensity of the primary emotions, as shown in Fig. 1 (c).
Building on this, EmoSphere-TTS [3] was previously introduced to model emotion style and intensity using a spherical emotion vector. However, the lack of consideration for emotion category distribution can lead to unnatural variations in certain styles and intensities. To address this, we propose an emotion-adaptive coordinate transformation that better models diverse emotional styles and intensities. Additionally, we introduce a joint attribute style encoder to enable emotion-controllable zero-shot TTS across a broader range of emotions, overcoming the limitations of relying on predefined emotion and speaker labels, which restrict flexibility in emotional expression. Furthermore, we achieve competitive performance solely through a conditional flow matching (CFM)-based decoder, eliminating the need for an additional discriminator module while enhancing emotional expressiveness and speech quality. The details of our approach are defined in the following subsections.
Figure 3: Training diagram of the EmoSphere++ framework. The framework consists of three main modules: the text encoder, the joint attribute style encoder, and the conditional flow matching (CFM) decoder. The right section illustrates the detailed structure of the joint attribute style encoder, which extracts global speaker, global emotion, and dimensional-driven emotion to form a joint attribute style embedding for emotional speech synthesis.